# Methods and stability tests associated with the sterile neutrino search using improved high-energy $\nu_{\mu}$ event reconstruction in IceCube

### v1

@@ Abstract
We provide supporting details for the search for a 3+1 sterile neutrino using data collected over eleven years at the IceCube Neutrino Observatory. The analysis uses atmospheric muon-flavored neutrinos from 0.5 to 100  TeV that traverse the Earth to reach the IceCube detector, and finds a best-fit point at $\sin^{2}(2\theta_{24})=0.16$ and $\Delta m^{2}_{41}=3.5$ eV2 with a goodness-of-fit p-value of 12% and consistency with the null hypothesis of no oscillations to sterile neutrinos with a p-value of 3.1%. Several improvements were made over past analyses, which are reviewed in this article, including upgrades to the reconstruction and the study of sources of systematic uncertainty. We provide details of the fit quality and discuss stability tests that split the data for separate samples, comparing results. We find that the fits are consistent between split data sets.
@@ I Introduction
Figure 1: Result of the IceCube search for $\nu_{\mu}$ disappearance with 11 years of data. The color scale shows the likelihood difference relative to the best-fit point. The best-fit point is marked with a black star, and the 90% (dotted), 95% (dashed), and 99% (solid) confidence level (C.L.) contours are drawn, assuming Wilks’ theorem with two degrees of freedom.
The IceCube Collaboration has published results from a $\nu_{\mu}$ disappearance search within a 3+1 model using eleven years of IceCube data in Ref. [7]. The analysis found a best-fit point at $\sin^{2}(2\theta_{24})=0.16$ and $\Delta m^{2}_{41}=3.5$ eV2 with a goodness-of-fit p-value of 12% and consistency with the no sterile neutrino (“null”) model given by a of p-value of 3.1%. Fig. 1 shows the result interpreted using Wilks’ theorem, assuming two degrees of freedom. Given this consistency with the null hypothesis, the result does not represent evidence for $\nu_{\mu}$ disappearance. This finding carries implications for the 3+1 neutrino landscape.
@@ II.1 Events in the IceCube Detector
Measurement of absorption and scattering of light in the ice is crucial to event reconstruction. We have found that at depths below 1450 m bubbles no longer play a significant role in the effective scattering (absorption) length of 400 nm light ranging from 20 m to 80 m (100 m to 400 m) depending on the depth. In situ LED light sources allow calibration of the light observed by DOMs showing that the bulk ice is tilted in a depth-dependent manner. This is incorporated in the ice model [16], which is used to model photon propagation for both simulation and reconstruction. Because the DOMs are inserted into a melted column of ice that re-freezes, the “hole ice” differs from the bulk ice, particularly due to bubbles (introduced during the hole-drilling process), and this effect is incorporated into the ice model.
@@ II.2.1 Atmospheric muon background rejection
In the previous analysis we achieved 99.91% purity of neutrino interaction events. In this analysis, we have taken muon rejection a step further to achieve similar purity while simultaneously increasing signal efficiency. The new step to mitigate the contamination from atmospheric muons implements a Boosted Decision Tree (BDT) using the AdaBoostClassifier algorithm [21]. The BDT was trained to differentiate between atmospheric muons and $\nu_{\mu}$ charged current interactions using 19 reconstructed observables (which were employed in the event selection of the previous analysis [20]). Fig. 2 demonstrates the effectiveness of the BDT in distinguishing atmospheric muons from muon neutrinos. Applying a BDT-score cut at 0.83 and using the reconstructed energy and zenith boundaries of the prior analyses [^1] The boundaries in the previous analysis were −$-1<\cos(\theta_{z}^{\mathrm{reco}})<0$ and $100.5<E_{\mathrm{reco}}/\mathrm{TeV}<$ 10, where $E_{\mathrm{reco}}$ was the likelihood-based estimator described in Sec. II.2.2., we achieved a nearly twofold increase in signal efficiency and a 50% reduction in atmospheric muon background.
@@ II.2.2 Energy estimation
The energy estimator used in this analysis uses the convolutional neural network (CNN) model architecture employed in previous analyses from IceCube [23, 24]. Charge and timing statistics of the pulses unfolded from the DOM waveforms are used as inputs to the CNN. The pulse time series is not directly used in the reconstruction, but rather, summary variables of the time distribution are used (i.e., overall charge, width of the pulse, time of first hit, etc.). The CNN model was trained on simulated $\nu_{\mu}$ CC interactions. The label it is trained on is the “visible energy” of each event, which is defined as follows. For muon neutrino interactions outside the instrumented volume, the visible energy is the muon’s energy when it enters the detector. For muon neutrino interactions occurring inside the instrumented volume, the visible energy is the muon’s energy plus the scaled energy of the hadronic shower [^2] The Cherenkov light yield for hadrons is extracted using light emission templates based on GEANT4 simulation [25]..
@@ II.2.3 Starting Event Identification
To discern between starting and through-going events, we employ a neural-network-based algorithm specifically designed to classify various event morphologies [26]. The distribution of starting track scores for events passing the BDT cut and within the energy-zenith boundaries of this analysis is illustrated in Fig. 4. Notably, $\nu_{\mu}$ charged current interactions within the instrumented volume, so-called starting events, cluster at high score values. In this analysis, we employ a classification threshold of 0.99 to distinguish between starting and through-going events. This choice results in 75% of the events in the data sample being identified as through-going and 25% as starting. The contamination of true through-going (starting) events in the starting (through-going) event sample is 0.2% (13%).
@@ III.3 Normalization
In our previous analysis, a 40% uncertainty was introduced in the normalization term for the conventional neutrino flux. Our updated treatment of this component now incorporates both shape and normalization uncertainties; therefore, we have revised the implementation of a normalization term. Key changes include extending the normalization term to cover non-conventional components and adjusting the prior width to 20%, to capture the influence of neutrino cross sections and muon energy losses in this analysis.
@@
III.4 Bulk ice
The presence of ice impurities, referred to as “dust,” between IceCube strings and the crystalline microstructure of the ice significantly influences the scattering and absorption of light. We address the uncertainty of dust concentration in different ice layers by leveraging LED flasher data [42]. To assess the impact of this effect, we conducted simulations using the SnowStorm method [43]. This method characterizes variations in the ice model through correlated amplitudes and phases of Fourier modes. The impact of these variations becomes negligible after the fourth mode in the reconstructed energy-zenith space. Fig. 8 illustrates that the primary impacts stem from the first amplitude and phase, while subsequent modes contribute second-order corrections. Consequently, our analysis incorporates five amplitude and four phase parameters as nuisance parameters, including a penalty term with correlated priors [43]. Additionally, we found that using simulation produced with the birefringent ice properties incorporated in the bulk ice model [44] has negligible impact on this analysis.
[…]
III.5 Local response of the DOMs
III.6 Neutrino attenuation
@@ III.7 Effect on sensitivity
Local response of DOMs (Sec. III.5)
@@ IV Fitting the 3+1 model
This analysis assumes that $\theta_{14}=\theta_{34}=0$ and leaves $\theta_{24}$ and $\Delta m^{2}_{41}$ as parameters in the fit. Fitting for more than two parameters is computationally costly, and so we begin this iteration of studies with only these two, recognizing that this is, in fact, a simplified 3+1 model.
@@
Resolution on $E_{\nu}$ and $\cos(\theta_{z}^{\nu})$ will smear the signature seen in the bottom oscillogram of Fig. 10. To see the effect, we compare the predictions for the best-fit and null models using the fitted nuisance parameters in each case. The bin-by-bin pull of the best fit minus the null prediction divided by the square root of the null prediction produces the prediction shown in the top panels of Fig. 11 for starting and through-going events. The apparent excess occurs because the normalization is allowed to float in the fit. One can see that the arc-like feature of the combined resonance and first oscillation maximum is retained even after detector effects are included. Detector smearing leads to an extended no-oscillation region below the arc. From 0.5 to 1 TeV, particularly in the case of starting events, an overall deficit from the fast oscillations can be seen at $\cos(\theta_{z}^{\mathrm{reco}})<-0.2$.
@@ IV.1 Frequentist Fit Results
Figure 12: Two-dimensional pulls. Pulls between data and best-fit expectations as a function of energy and zenith.
@@
The bottom panels in Fig. 11 compare the bin-wise pulls between the data and both the best-fit and null expectations. Purple (orange) bins indicate better agreement with the best-fit prediction (null hypothesis). Across all bins, differences remain within the ±$\pm 1\sigma$ range, with the largest deviations of ±$\pm 0.6\sigma$, primarily observed in the most horizontal events within the through-going sample—a point discussed further in Sec. V. Up to statistical fluctuations, the areas showing better agreement with the best-fit prediction coincide with regions where the pulls between the best-fit and null hypothesis are more prominent, as shown in the top panels in Fig. 11.
@@ IV.2 Bayesean Fit Results
To complement the frequenstist result, we have also developed a Bayesian analysis. In this approach, the model evidence for each sterile neutrino hypothesis is computed by integrating the likelihood over the nuisance parameters using MultiNest [49]. In order to avoid dependence on the physics parameter priors, we compare each sterile point in our ($\sin^{2}(2\theta_{24}),\Delta m^{2}_{41}$) grid to the null hypothesis. We use the same priors as in the frequentist analysis for the nuisance parameters.
@@ V Post-unblinding stability tests
Figure 17 (top left) compares the results if starting and through-going events are fit separately compared to the simultaneous fit used in the final result. The larger through-going event sample (red), with 274,309 events, retains a 95% C.L. contour that overlaps the final fit well and has a best fit contained within the final sample 85% C.L. allowed region. The smaller starting sample (blue), with 93,762 events, does not have a 95% C.L. closed contour. However, the best fit agrees with the final allowed region. Also, the 95% C.L. limit contour follows the final best-fit region, as expected.
Figure 17 (top middle) divides the sample into low- ($<10$ TeV) and high-energy ($>10$ TeV) events. We chose 10 TeV as the point to split the data sets because this is the upper bin in the previous analysis [17]. The samples have 361,633 and 6,438 events, respectively. Although neither result shows a closed contour, both results have best-fit points within the final sample allowed region and contours that behave as expected. Hence, we conclude that there is no evidence of bias in the analysis due to the description of the fluxes or the energy reconstruction.
Figure 17 (top right) splits the sample in the azimuthal direction, which is the angle in a horizontal plane perpendicular to the IceCube strings. This test examines bias in the reconstruction arising from ice tilt, which is particularly interesting given recent measurements indicating not only tilt along the direction orthogonal to the ice flow but also new components not accounted for in the simulation used in this analysis [51]. Six angles are chosen to reflect three known tilts, where three are parallel to the ice sheet and three are perpendicular. This substantially reduces the statistics, with approximately 60,000 events in each sample. The behavior of the contours and the best-fit points is as expected. The best-fit points scatter more than for the other cases due to the lower statistics. This test is quite important for an analysis that searches for energy dependence associated with an oscillation and a resonance, but there is no evidence of disagreement between the split data sets and the final result.
Figure 17: Split sample fits. Each plot corresponds to a reconstructed variable that has been divided into different regions. Each color line shows the 95% C.L. in a given range, and the stars represent the best-fit point. The grey region shows the final result of this analysis, and the black circle is the best-fit point. Top panels: starting vs. through-going (left), low vs. high energy (middle), and for six azimuthal angles (right). Bottom panels: analysis performed with sample split at different zenith angles, with the more up-going sample results in the left plot, and their more horizontal counterparts in the right plot.
@@
For this cross-check, we split the data into increasing samples from the edges of the analysis cuts. The results are shown in the bottom panels of Fig. 17. In the left (right) plot, more horizontal (vertical) events are progressively excluded from the fit. We observe that the contours shrink as more horizontal events are removed from the analysis, and the fits involving these horizontal events prefer a slightly higher $\Delta m_{41}^{2}$. Nonetheless, the results remain consistent with those drawn from the full-zenith fit.

### v2

@@ Abstract
We provide supporting details for the search for a 3+1 sterile neutrino using data collected over 10.7 years at the IceCube Neutrino Observatory. The analysis uses atmospheric muon-flavored neutrinos from 0.5 to 100  TeV that traverse the Earth to reach the IceCube detector, and finds a best-fit point at $\sin^{2}(2\theta_{24})=0.16$ and $\Delta m^{2}_{41}=3.5$ eV2 with a goodness-of-fit p-value of 12% and consistency with the null hypothesis of no oscillations to sterile neutrinos with a p-value of 3.1%. Several improvements were made over past analyses, which are reviewed in this article, including upgrades to the reconstruction and the study of sources of systematic uncertainty. We provide details of the fit quality and discuss stability tests that split the data for separate samples, comparing results. We find that the fits are consistent between split data sets.
@@ I Introduction
Figure 1: Result of the IceCube search for $\nu_{\mu}$ disappearance with 10.7 years of data. The color scale shows the likelihood difference relative to the best-fit point. The best-fit point is marked with a black star, and the 90% (dotted), 95% (dashed), and 99% (solid) confidence level (C.L.) contours are drawn, assuming Wilks’ theorem with two degrees of freedom.
The IceCube Collaboration has published results from a $\nu_{\mu}$ disappearance search within a 3+1 model using 10.7 years of IceCube data in Ref. [7]. The analysis found a best-fit point at $\sin^{2}(2\theta_{24})=0.16$ and $\Delta m^{2}_{41}=3.5$ eV2 with a goodness-of-fit p-value of 12% and consistency with the no sterile neutrino (“null”) model given by a of p-value of 3.1%. Fig. 1 shows the result interpreted using Wilks’ theorem, assuming two degrees of freedom. Given this consistency with the null hypothesis, the result does not represent evidence for $\nu_{\mu}$ disappearance, but it provides valuable information for our understanding of the 3+1 neutrino landscape.
@@ II.1 Events in the IceCube Detector
Measurement of absorption and scattering of light in the ice is crucial to event reconstruction. At depths below 1450 m, bubbles no longer play a significant role in the effective scattering or absorption lengths of 400 nm light. Depending on the depth, the effective scattering length ranges from 20 m to 80 m, and the absorption length ranges from 100 m to 400 m. In situ LED light sources allow calibration of the light observed by DOMs showing that layers of glaciologically contemporaneous ice are not perfectly horizontal but vary in depth across the array, usually referred to as bulk ice tilt. This is incorporated in the ice model [16], which is used to model photon propagation for both simulation and reconstruction. Because the DOMs are inserted into a melted column of ice that re-freezes, the “hole ice” differs from the bulk ice, particularly due to bubbles (introduced during the hole-drilling process), and this effect is incorporated into the ice model.
@@ II.2.1 Atmospheric muon background rejection
In the previous analysis, we achieved 99.91% purity of $\nu_{\mu}$ CC events. In this analysis, we have taken muon rejection a step further to achieve similar purity while simultaneously increasing signal efficiency. The new step to mitigate the contamination from atmospheric muons implements a Boosted Decision Tree (BDT) using the AdaBoostClassifier algorithm [21]. The BDT was trained to differentiate between atmospheric muons and $\nu_{\mu}$ charged current interactions using 19 reconstructed observables (which were employed in the event selection of the previous analysis [20]). Although the BDT was trained under the null hypothesis of no sterile neutrinos, we found that the BDT score distribution remained stable across different sterile neutrino hypotheses. Fig. 2 demonstrates the effectiveness of the BDT in distinguishing atmospheric muons from muon neutrinos. Applying a BDT-score cut at 0.83 and using the reconstructed energy and zenith boundaries of the prior analyses [^1] The boundaries in the previous analysis were −$-1<\cos(\theta_{z}^{\mathrm{reco}})<0$ and $100.5<E_{\mathrm{reco}}/\mathrm{TeV}<$ 10, where $E_{\mathrm{reco}}$ was the likelihood-based estimator described in Sec. II.2.2., we achieved a nearly twofold increase in signal efficiency and a 50% reduction in atmospheric muon background, reaching a 99.93% purity of $\nu_{\mu}$ CC events. The agreement between data and simulation within the selected region supports the estimated purity.
@@ II.2.2 Energy estimation
The energy estimator used in this analysis uses the convolutional neural network (CNN) model architecture employed in previous analyses from IceCube [23, 24]. Charge and timing statistics of the pulses unfolded from the DOM waveforms are used as inputs to the CNN. The pulse time series is not directly used in the reconstruction, but rather, summary variables of the time distribution are used (i.e., overall charge, width of the pulse, time of first hit, etc.). The CNN model was trained on simulated $\nu_{\mu}$ CC interactions. The label it is trained on is the “visible energy” of each event, which is defined as follows. For muon neutrino interactions outside the instrumented volume, the visible energy is the muon’s energy when it enters the detector. For muon neutrino interactions occurring inside the instrumented volume, the visible energy is the muon’s energy plus the scaled energy of the hadronic shower [^2] The Cherenkov light yield for hadrons is extracted using light emission templates based on GEANT4 simulation [25].. Therefore, the visible energy closely matches the true neutrino energy for starting events.
@@ II.2.3 Starting Event Identification
To discern between starting and through-going events, we employ a neural-network-based algorithm specifically designed to classify various event morphologies [26]. The distribution of starting track scores for events passing the BDT cut and within the energy-zenith boundaries of this analysis is illustrated in Fig. 4. Notably, $\nu_{\mu}$ charged current interactions within the instrumented volume, so-called starting events, cluster at high score values. In this analysis, we employ a classification threshold of 0.99 to distinguish between starting and through-going events. This threshold was chosen to obtain a very pure sample of starting events, ensuring good energy resolution within that sample. This choice results in 75% of the events in the data sample being identified as through-going and 25% as starting. The contamination of true through-going events in the starting event sample is 0.2%. Conversely, the contamination of true starting events in the through-going event sample is 13%.
@@ III.1 Conventional flux
The main difference compared to the previous parametrization is observed in the $\nu_{\mu}/\bar{\nu}_{\mu}$ ratio, where DAEMONFLUX shows larger $\nu_{\mu}$ contributions above 100 GeV and smaller uncertainties. Additionally, in the TeV regime, DAEMONFLUX and the previous model predict similar $\nu_{\mu}+\bar{\nu}_{\mu}$ fluxes (more details about these comparisons can be found in Ref. [13]).
We computed the sensitivity to a sterile neutrino using both models, as well as performed mismodeling tests by injecting and fitting with different parametrizations. In both cases, we observed that the most conservative result was obtained when fitting with DAEMONFLUX. The primary reason is that DAEMONFLUX has more flexibility in changing the shape of the spectrum, which is critical when looking for sterile-like distortions in the flux.
@@ III.3 Normalization
In our previous analysis, a 40% uncertainty was introduced in the normalization term for the conventional neutrino flux. Our updated treatment of this component now incorporates both shape and normalization uncertainties; therefore, we have revised the implementation of a normalization term. Key changes include extending the normalization term to cover non-conventional components and adjusting the prior width to capture the influence of neutrino cross sections and muon energy losses in this analysis.
@@
In conclusion, the prior width of the normalization term has been updated to 20%.
III.4 Neutrino attenuation
@@
III.5 Bulk ice
The presence of ice impurities, referred to as “dust,” between IceCube strings and the crystalline microstructure of the ice significantly influences the scattering and absorption of light. We address the uncertainty of dust concentration in different ice layers by leveraging LED flasher data [45]. To assess the impact of this effect, we conducted simulations using the SnowStorm method [46]. This method characterizes variations in the ice model through correlated amplitudes and phases of Fourier modes. The impact of these variations becomes negligible after the fourth mode in the reconstructed energy-zenith space.
As illustrated in Fig. 8, the primary impacts stem from the zeroth-mode amplitude and the first-mode phase (i.e., Amp. 0 and Phs. 1), while subsequent modes contribute second-order corrections. The Amp. 0 sets the absolute scale of absorption and scattering in the detector, and the Phs. 1 encodes the most relevant depth-dependent features. The latter affects starting and through-going events differently. The main reason is that the light pattern is partially contained within the detector for starting events, making them more sensitive to depth-dependent ice properties.
[…]
Consequently, our analysis incorporates five amplitude and four phase parameters as nuisance parameters, including a penalty term with correlated priors as shown in Fig. 9. Additionally, we found that using simulation produced with the birefringent ice properties incorporated in the bulk ice model [47] has negligible impact on this analysis.
Figure 9: Correlation for bulk ice parameters. Correlation matrix of the bulk ice parameters used in this analysis extracted from Ref. [46].
III.6 Local response of the DOMs
@@ III.7 Effect on sensitivity
Local response of DOMs (Sec. III.6)
@@ IV Fitting the 3+1 model
This analysis assumes that $\theta_{14}=\theta_{34}=0$ and leaves $\theta_{24}$ and $\Delta m^{2}_{41}$ as parameters in the fit. Fitting for more than two parameters is computationally costly, and so we begin this iteration of studies with only these two, recognizing that this is, in fact, a simplified 3+1 model. The impact of $\theta_{14}$ is negligible in this analysis since the $\nu_{e}/\nu_{\mu}$ ratio is higher than 20 in the TeV regime [13], whereas setting $\theta_{34}=0$ yields conservative constraints on $\theta_{24}$ [49].
@@
Resolution on $E_{\nu}$ and $\cos(\theta_{z}^{\nu})$ will smear the signature seen in the bottom oscillogram of Fig. 11. To see the effect, we compare the predictions for the best-fit and null models using the fitted nuisance parameters in each case. The bin-by-bin pull of the best fit minus the null prediction divided by the square root of the null prediction produces the prediction shown in the top panels of Fig. 12 for starting and through-going events. The apparent excess occurs because the normalization is allowed to float in the fit. As indicated in Table 2, the normalization for the best-fit hypothesis is higher than that for the null hypothesis. This difference arises because the fast oscillations reduce the expected flux below the resonance, as illustrated in Fig.11. In addition, one can see that the arc-like feature of the combined resonance and first oscillation maximum is retained even after detector effects are included. Detector smearing leads to an extended no-oscillation region below the arc. From 0.5 to 1 TeV, particularly in the case of starting events, an overall deficit from the fast oscillations can be seen at $\cos(\theta_{z}^{\mathrm{reco}})<-0.2$.
@@ IV.1 Frequentist Fit Results
Figure 13: Two- and one-dimensional pulls. Left panels: pulls between data and best-fit expectations as a function of energy and zenith. Right panel: One-dimensional distribution of the pulls from the 880 bins used in this analysis. The orange line represents the Gaussian fit.
@@
The bottom panels in Fig. 12 compare the bin-wise pulls between the data and both the best-fit and null expectations. Purple (orange) bins indicate better agreement with the best-fit prediction (null hypothesis). Across all bins, differences remain within the ±$\pm 1\sigma$ range, with the largest deviations of ±$\pm 0.6\sigma$, primarily observed in the most horizontal events within the through-going sample—a point discussed further in Sec. V. Up to statistical fluctuations, the areas showing better agreement with the best-fit prediction coincide with regions where the pulls between the best-fit and null hypothesis are more prominent, as shown in the top panels in Fig. 12. The main feature of the data that leads the fit to prefer the best-fit over the null hypothesis appears in the through-going sample for $\cos(\theta_{z}^{\textrm{reco}})<-0.4$ at energies around 10 TeV. In fact, we observe that the null rejection increases when studying these regions separately, as discussed in Sec. V.
@@ IV.2 Bayesean Fit Results
To complement the frequenstist result, we have also developed a Bayesian analysis. We have followed the same method as in the previous analysis [20], which avoids dependence on the physics parameter priors as described in Ref. [50]. In this approach, the model evidence is computed for each sterile neutrino hypothesis by integrating the likelihood over the nuisance parameters using MultiNest [51]. Then, we compare each sterile point in our ($\sin^{2}(2\theta_{24}),\Delta m^{2}_{41}$) grid to the null hypothesis. We use the same priors as in the frequentist analysis for the nuisance parameters.
@@ V Post-unblinding stability tests
Figure 18 (first panel) compares the results if starting and through-going events are fit separately compared to the simultaneous fit used in the final result. The larger through-going event sample (red), with 274,309 events, retains a 95% C.L. contour that overlaps the final fit well and has a best fit contained within the final sample 95% C.L. allowed region. The smaller starting sample (blue), with 93,762 events, does not have a 95% C.L. closed contour. However, the best fit agrees with the final allowed region. Also, the 95% C.L. limit contour follows the final best-fit region, as expected.
Figure 18 (second panel) divides the sample into low- ($<10$ TeV) and high-energy ($>10$ TeV) events. We chose 10 TeV as the point to split the data sets because this is the upper bin in the previous analysis [17]. The samples have 361,633 and 6,438 events, respectively. Although neither result shows a closed contour, both results have best-fit points within the final sample allowed region and contours that behave as expected. Hence, we conclude that there is no evidence of bias in the analysis due to the description of the fluxes or the energy reconstruction.
Figure 18 (third panel) splits the sample in the azimuthal direction, which is the angle in a horizontal plane perpendicular to the IceCube strings. This test examines bias in the reconstruction arising from ice tilt, which is particularly interesting given recent measurements indicating not only tilt along the direction orthogonal to the ice flow but also new components not accounted for in the simulation used in this analysis [53]. Six angles are chosen to reflect three known tilts, where three are parallel to the ice sheet and three are perpendicular. This substantially reduces the statistics, with approximately 60,000 events in each sample. The behavior of the contours and the best-fit points is as expected. The best-fit points scatter more than for the other cases due to the lower statistics. This test is quite important for an analysis that searches for energy dependence associated with an oscillation and a resonance, but there is no evidence of disagreement between the split data sets and the final result.
Figure 18: Split sample fits. Each plot corresponds to a reconstructed variable that has been divided into different regions. Each color line shows the 95% C.L. in a given range, and the stars represent the best-fit point. The grey region shows the final result of this analysis, and the black circle is the best-fit point. From left to right: starting vs. through-going, low vs. high energy, six azimuthal angles, and four zenith angles (note: the island observed in the contour for −$-1<\cos\theta<-0.75$ indicates exclusion at 95% C.L.).
@@
For this cross-check, we split the data into different zenith directions. The results are shown in the fourth panel of Fig. 18. We observe that none of the zenith samples show a closed contour at 95% C.L. However, the best-fit points remain in the same region as those drawn from the full-zenith fit.
Figure 19: Comparison with other experiments. The 90% and 99% C.L. contours (red lines) compared to world data at 90% CL. Left panel: long-baseline experiments using $\nu_{\mu}$-mode beam (MiniBooNE-SciBooNE [55], MINOS [56], CDHS and CCFR [1]). Center panel: long-baseline experiments using $\bar{\nu}_{\mu}$-mode beam (MiniBooNE-SciBooNE [57]). Right panel: atmospheric neutrino experiments (SuperK [58] and DeepCore [59, 60]).
@@
Finally, Fig. 19 presents this result in comparison with previous measurements of $\nu_{\mu}$ disappearance from other experiments. The 90% C.L. allowed region from this result indicates an increased tension with the constraints from long-baseline experiments, namely MINOS and MiniBooNE-SciBooNE. The tension is smaller with experiments using atmospheric neutrinos.